solve the absolute value equation. express numbers as integers or simplified fractions.
|(5b)/2 + 7/2| - 15 = -4

Answer by Boreal(15235)

abs((5b/2)+3.5)=11
therefore
2.5b+3.5=11, so 2.5b=7.5 and b=3
and
2.5b+3.5=-11, so 2.5 b=-14.5 and b=-5.8 (-29/5)
abs (7.5+3.5)=11 and abs (-14.5+3.5)=11